OpenStudy (anonymous):

an = -10 – 3(n – 1) I need the next 5 terms and the 10th term

OpenStudy (anonymous):

The most basic way to do this is to plug 1, 2, 3, 4, 5, and 10 into the equation for n

OpenStudy (anonymous):

For example, a1 = -10 - 3(1-1) = -10 -3(0) = -10